# In situ coherent X-ray scattering reveals polycrystalline structure and discrete annealing events in strongly-coupled nanocrystal superlattices

### 11 Jan 2024 v1

@@ I Introduction
Here, we capture a snapshot of an evolving SL formed from electrostatically stabilized NCs and also observe its annealing events to elucidate their nature. To achieve stability in solution for Bragg diffraction and CDI reconstruction, we developed a strategy to immobilize SLs in capillaries filled with quartz wool. Performing Bragg CDI[34, 35, 36, 37] on the fcc (111) SL peak, we resolve a 600 nm core of apparent high electron density, surrounded by a few-micron apparently lower electron density region. We attribute the lower densities to crystallites that are not well-aligned with the measured (111) peak. Moreover, in using XPCS on the same SL, we are able to both validate the sufficient degree of immobilization in solution and elucidate the nature of SL annealing. Taken together, we conclude that SLs in solution form via a fast nucleation and growth step resulting in structures with a high density of grain boundaries. Subsequent annealing of these grain boundaries via discrete shearing events enables further ordering of the SLs.
@@ II Results
To perform the experiments, SLs of 4.6-nm diameter Au NCs capped with thiostannate () ligands are prepared by adding 75 mM salt to an electrostatically stabilized colloidal solution in hydrazine (), as described previously[1, 13] (Fig. 1(a)). A typical small-angle X-ray scattering (SAXS) intensity pattern, , as a function of momentum transfer, , of these samples measured previously is shown in Fig. 1(b). Approximately 24 hours after the SLs are formed, they are injected into a thin-walled borosilicate capillary into which quartz wool has been inserted to immobilize them while maintaining their solvent environment to enable further evolution (Fig. 1(c), S1). The capillaries are securely sealed while still in an inert environment, and are mounted and moved into the center of rotation at the X-ray correlation spectroscopy (XCS) endstation of the Linac Coherent Light Source (LCLS) [38] (Fig. 1(d)). The experiment was carried out at 11.8 keV, just below the Au absorption edge, using the Si(111) monochromator in a SAXS geometry with a sample-to-detector distance of 7.7 m, measuring to 0.15 in order to capture the (111) SL peak as shown in Fig. 1(e). The X-ray beam, consisting of 50 fs pulses at 120 Hz, was focused with low divergence to m and attenuated to 5 photon/s to preserve the integrity of the sample and its solution environment. To characterize aspects of the data we commonly refer below to the location in momentum space on the detector relative to the direct beam () and also using a reduced scattering vector , defined as , where is the (111) reciprocal lattice vector.
@@
2(b) presents the apparent three-dimensional real space electron density, replotted in an orthonormal coordinate basis (see supplement, Fig. S2). Each (cubic) voxel has a side length of 0.071 m. The real space resolution is given by divided by the dimensions of the sampled reciprocal space region and is then nearest-neighbor interpolated to obtain isotropic sampling. The corresponding reconstructed diffraction pattern, obtained via Fourier transform of the electron density, is shown in 2(c). Comparison between 2(a) and 2(c) shows good agreement between the measured and reconstructed diffraction patterns, especially in brighter regions where the signal-to-noise ratio in the data is higher. In particular, the reconstructed pattern preserves the fringe spacing and overall contours of the data. It also features singularities (dark spots) in the already low-intensity regions of the diffraction pattern that are not clearly visible in the measured detector image. These characteristics persist across reconstructions with different random initial starting points. The apparent density of the reconstruction in 2(b) features a higher-density central region surrounded by a few-micron shell over which the density tapers off by more than an order of magnitude (Fig. S3). In an effort to understand whether the lower-magnitude outer regions seen in pale blue in 2(b) are physical or simply a reconstruction artifact, we perform a set of reconstructions with various support sizes in which the support is unaltered during the reconstruction algorithm. 2(d) compares a line cut of the data (black circles) to line cuts of the Fourier transform of the reconstructed density as a function of support diameter. The reconstructed reciprocal space intensities reproduce the features of the data well only for sufficiently large supports that extend well beyond the bounds of the higher-magnitude central region of the density. This analysis therefore suggests that there is indeed a physical basis for the lower-magnitude region featured in 2(b). Furthermore, we obtain similar results from data collected on a different SL in the same capillary (see Fig. S4, S5 and accompanying text). We defer proposing origins of this apparent lower-magnitude density region to the Discussion in order to first present the XPCS results for this same SL structure.
@@
To characterize aspects of SL annealing we performed additional analysis on the dataset shown in Fig. 3. To elucidate the scale-dependent dynamics within the SL, we calculate correlation functions as a function of distance from the Bragg peak, as shown in Fig. 4(a). We divide the detector ROI containing the peak into annular regions bounded by contours of constant , and we further subdivide each region along the longitudinal and transverse components of , marked in gold and blue, respectively. In this SAXS geometry, the longitudinal direction is parallel to , and the transverse direction is tangent to the (111) diffraction ring. For each annular region, we fit the corresponding autocorrelation function, , to a compressed/stretched exponential of the form , where is a baseline, is the speckle contrast, is the decorrelation rate, and is a stretching exponent. Fig. 3(b)-(c) shows example autocorrelations and fits for respective longitudinal and transverse directions at (see also Figs. S6, S7, S8 for data and fits at additional values). Such stretched/compressed exponentials often appear in systems with heterogeneous dynamics, where the resulting distribution of time scales leads to distortion of the simple exponential decay or Gaussian correlation forms [43]. Note that a speckle contrast higher than 1 arises when there are large variations in intensity (e.g., diffraction fringes) within the ROI. Because this analysis depends on the choice of center for the annular regions of interest and there is ambiguity in how this choice should be made, we repeat the analysis for many different choices of center, using pixels within a circular region (indicated in red in Fig. 3(a)) with a 5 pixel radius around the intensity center of mass, and averaging the resultant fit parameters together. Fig. 3(d) shows the decorrelation rate as a function of for each of the radial and transverse directions, respectively. The corresponding stretching exponents, , are shown in Fig. 3(e). The behavior of at low (region i, in gray) is complex and non-monotonic. The dispersion contrasts with that of, for instance, a simple, freely-diffusing colloidal system, where [26]. Unlike a diffusive system, which has no characteristic length scale, however, our system has a clear length scale of relevance, namely the size of the SL. For values around this value we should not expect a simple dispersion. Only for sufficiently small length scales (region ii, white) do we recover standard scattering behavior of faster decorrelation at higher . In the transverse direction, the decorrelation rate is on the order of Hz and increases with increasing , while in the longitudinal direction we do not resolve such dispersive behavior (and cannot reliably extract fits at the highest values). This trend in decorrelation rate in the transverse direction, and the lack thereof in the longitudinal direction, persist even if one fits to a fixed exponent of instead (Fig. S9). In order to investigate the source of this wavevector-dependent time scale, we calculate a two-time correlation function in Fig. 3(f) using the ring-shaped region of Fig. 3(a) around Two-time analysis enables us to observe from when in the data acquisition period the time scale arises and whether it is irreversible [44, 45, 46]. The raw data were first binned down to 0.1 s intervals to more readily visualize larger jumps in intensity. We observe distinct square-like patterns in the two-time correlation, corresponding to sudden shifts in the diffraction intensity at 17 s and 90 s.
@@ III Discussion
This work extends the reach of coherent diffractive imaging by probing a solid, immobilized structure in its native liquid environment[47]. To be able to extract meaningful conclusions about SL structure and evolution from our diffraction data via CDI and XPCS analysis techniques, the SLs must be sufficiently immobilized in the X-ray beam. The low- behavior of the SL Bragg peak is governed by long-range order in the SL; thus the oscillations that we observe in the autocorrelation functions in 2, which feature prominently at low , are likely produced by a bulk motion of the SL, such as small librations of the SL in solution. Perturbed by solvated particles or external vibrations, the SLs sway periodically on the quartz wool, producing intensity fluctuations at a rate much slower than the X-ray pulse repetition rate of 120 Hz. While the SLs are therefore not perfectly immobilized in the capillary, these perturbations are quite small. Some challenges to extracting the SL motion from the data directly include: the time-resolved data is a 2D cross-section of the 3D diffraction pattern; rotations do not easily map onto a rectilinearly-sampled dataset, and it is not straightforward to determine intensity fluctuations from an intensity autocorrelation. We can, however, attempt to place limits on the extent of the motion using the resolution of the measured static 3D diffraction pattern. In particular, we identify sharp peaks in individual pixel intensities as a function of rocking angle, with a full-width at half-maximum of five angular steps, or 0.078 (Fig. S10). For the in-plane direction, we see fringes, normal to the transverse direction, with a full-width at half-maximum of roughly five pixels (Fig. S11). A rotation of the SL of only approximately 0.25 about the direct beam would shift a feature in the diffraction pattern by five pixels in the transverse direction. Since the data acquisition period (2 s) for each rocking angle is roughly an order of magnitude larger than the autocorrelation oscillation period, these fringe widths serve as upper bounds on the extent to which our intensity dataset may be blurred by SL librations. That is, we expect the SL rotation about the rocking axis to be no more than about 0.08 (a maximum motion of a 4 nm arc length for a 3 m structure) and the rotation about the direct beam to be no more than about 0.25 (a maximum motion of a 13 nm arc length for a 3 m structure). We cannot easily quantify rotations about a third axis orthogonal to these first two, but since there is no reason to expect that the superstructure would preferentially oscillate about such an axis, we assume a similar value for it as well. Therefore, while we cannot quantify the SL motion exactly, that we are able to resolve such sharp features in our dataset is an encouraging sign that the sample is sufficiently immobilized for the purposes of XPCS and CDI analyses.
@@
We now turn to a discussion of the long-time behavior of the SLs, as revealed by our XPCS results. We first observe that the dispersive increase (Fig. 3(d)) at high in the transverse fit parameters extracted from the XPCS data reflects conventional dynamical behavior in which scattering decorrelates more quickly at higher wavevector () [26]. Moreover, wavevectors in the transverse direction are sensitive to shear-like motions of the SL, analogous to the sensitivity to transverse acoustic phonons near Bragg peaks[48, 49]. Therefore, since we observe this dispersion along the transverse direction but not along the longitudinal direction, we interpret the time scale as that of shear-like fluctuations of the SL. We see, furthermore, that this one-time decorrelation is accompanied by square-like patterns in the two-time correlation (Fig. 3(f)), a phenomenon that we also observe in an additional experimental dataset (Fig. S12). In a third experimental dataset (Fig. S13, S14), we do not observe any decay in the autocorrelation functions at long delay, and the corresponding two-time correlation function (Fig. S15) is highly uniform and featureless, displaying no such rapid discontinuities as those seen in Fig. 4(f). Since the absence or presence of a one-time decorrelation rate is thus accompanied by the absence or presence of these discontinuities in the two-time correlation, the decorrelation is likely dominated by sudden changes in the diffraction pattern instead of small fluctuations therein. While we cannot rule out the possibility that these sudden changes in the diffraction pattern are merely due to bulk translations or rotations of the SL in solution, the increasing dispersion in decorrelation rate at high is indicative of internal SL dynamics. Furthermore, from prior ensemble SAXS measurements on NC SLs [13], we expect SLs to continue to evolve (via growth and defect annealing) even many hours after they are initially nucleated, provided their solution phase environment is maintained. Taken together, these features in the data suggest that the high- fluctuations in the SL Bragg peak are driven by sudden shear-like “avalanches”[50, 51] of NCs rather than a continuous annealing process. Such events are roughly analogous to the microstructural avalanches previously observed in martensitic metallic alloys [46, 45]. In our previous in situ SAXS experiments of SL formation, SLs were shown to evolve at different rates depending on the quench depth at which they were formed. At shallow quenches, highly ordered SLs formed but did not evolve much after their initial formation. In contrast, at slightly deeper quenches, more disordered SLs form initially but they continue to evolve over the course of the several hour observation period[13]. This observation is consistent with a picture in which annealing proceeds via such shear-like avalanches, in which these strain-relieving events are more likely to occur in regions of high disorder[52, 53, 54]. Moreover, previous ensemble measurements also find that the SL Bragg peak width evolves according to a power law over time, a relationship we would expect to arise from the occurrence of avalanches within many SLs within an ensemble throughout a bulk measurement window[55].
@@
Returning to the real space reconstruction obtained from CDI presented in 2 we consider interpretations of the decrease in electron density that arises moving outward from the center of the SL. One possibility is that the reconstruction algorithm has failed, assigning low non-zero values to regions of real space that extend beyond the boundary of the object. Another possibility is that the diffraction data is blurred due to the SL librations discussed above, thus yielding artifacts on large length scales in the reconstruction. A further alternative is that inhomogeneities in the beam path (due to the presence of solvent, quartz wool, and capillary glass) induce an effective decoherence in the beam and distort the diffraction signal. There are several reasons, however, to believe that the low-density regions are not an experimental artifact. First, the analysis of size-dependence presented in 2(d)suggests that there is diffracting material over a span of 2-3 m, extending beyond the high density region of the reconstruction. Additionally, we observe similar such low-density regions in the second CDI dataset that we analyzed (Figure S5). For this latter dataset the reconstructed object, including the low-density region, is smaller, in accordance with a larger fringe spacing. If beam contamination were the driving force behind a reconstruction artifact, we would expect to see its impact appear on similar length scales across all datasets. Furthermore, we can estimate an upper bound for the magnitude of phase distortion in the scattered beam. For quartz wool fibers of 9-30 m, supposing that the part of the beam passes through at worst the edge of one fiber, we would expect a relative phase shift of around . This places an upper bound on the phase distortion due to the quartz wool, and is small enough to preserve a good CDI signal (see supplemental material for quantitative estimates). One physical interpretation of these apparent lower electron densities is that they are produced by regions of greater disorder in the SL. By analyzing and comparing hypothetical ordered and disordered SL structures, we attempt to assess the impact of at least one kind of defect on such a Bragg CDI reconstruction.
@@
The combination of CDI and XPCS analysis together has proven to be a valuable means to obtain a coherent description of a self-assembled SL structure and its evolution in solution that could not otherwise have been realized. We previously found via in situ SAXS that initial SL nucleation takes place on a rapid (sub-second) time scale, and then growth and annealing progress over the course of minutes and hours. Taking all of our findings together, one interpretation of the apparent density profile we observe from CDI is that the SL crystallinity decays away from its center, with regions of greater disorder towards the surface. We would not expect such a disordered shell, however, to be so large in proportion to the crystalline core. In addition, real space optical and electron microscopy images of this system following a quench show large, micron sized polycrystalline structures (Fig. S1). Therefore a more likely explanation is that the SL (111) peak analyzed in 2 is a single crystal grain of a larger cluster of aggregated SLs, each of which nucleated separately. Our analysis of hypothetical polycrystals in 5 demonstrates that in performing CDI on a polycrystalline structure whose other grains do not meet the Bragg condition, we would expect to see regions of apparent low density in the reconstructed object. This polycrystalline picture also accords well with the account of sub-lattice dynamics indicated by correlation analysis. In particular, the presence of defects such as slip planes or grain boundaries between neighboring crystallites can lead to a build-up of strain in the SL. Once this strain becomes sufficiently large, annealing should occur via discrete, large-scale shear motions. Based on the data analyzed here and in previous work, we thus propose that Au NC SLs quickly nucleate and then coalesce into polycrystalline structures, with shear annealing occurring within and/or between crystallites via strain-relieving avalanches.

### 03 Mar 2024 v2

@@ I Introduction
Here, we capture a snapshot of an evolving SL formed from electrostatically stabilized NCs and also observe its annealing events to elucidate their nature. To achieve stability in solution for Bragg diffraction and CDI reconstruction, we developed a strategy to immobilize SLs in capillaries filled with quartz wool. Performing Bragg CDI [34, 35, 36, 37] on the fcc (111) SL peak, we resolve a 600 nm core of apparent high electron density, surrounded by a few-micron apparently lower electron density region. We attribute the lower densities to crystallites that are not well-aligned with the measured (111) peak. Moreover, in using XPCS on the same SL, we are able to both validate the sufficient degree of immobilization in solution and elucidate the nature of SL annealing. Combining these findings, we conclude that SLs in solution form via a fast nucleation and growth step resulting in structures with a high density of grain boundaries. Subsequent annealing of these grain boundaries via discrete shearing events enables further ordering of the SLs.
@@ II Results
To perform the experiments, SLs of 4.6-nm diameter Au NCs capped with thiostannate () ligands are prepared by adding 75 mM salt to an electrostatically stabilized colloidal solution in hydrazine (), as described previously [1, 13] (Fig. 1(a)). A typical small-angle X-ray scattering (SAXS) intensity pattern, , as a function of momentum transfer, , of these samples measured previously is shown in Fig. 1(b). Approximately 24 hours after the SLs are formed, they are injected into a thin-walled borosilicate capillary into which quartz wool has been inserted to immobilize them while maintaining their solvent environment to enable further evolution (Fig. 1(c), Fig. 6). The capillaries are securely sealed while still in an inert environment, and are mounted and moved into the center of rotation at the X-ray correlation spectroscopy (XCS) endstation of the Linac Coherent Light Source (LCLS) [38] (Fig. 1(d)). The experiment was carried out at 11.8 keV, just below the Au absorption edge, using the Si(111) monochromator in a SAXS geometry with a sample-to-detector distance of 7.7 m, measuring to 0.15 in order to capture the (111) SL peak as shown in Fig. 1(e). The X-ray beam, consisting of 50 fs pulses at 120 Hz, was focused with low divergence to m and attenuated to 5 photon/s to preserve the integrity of the sample and its solution environment. To characterize aspects of the data we commonly refer below to the location in momentum space on the detector relative to the direct beam () and also using a reduced scattering vector , defined as , where is the (111) reciprocal lattice vector.
@@
Fig. 2(b) presents the apparent three-dimensional real space electron density, replotted in an orthonormal coordinate basis (see Fig. 7). Each (cubic) voxel has a side length of 0.071 m. The real space resolution is given by divided by the dimensions of the sampled reciprocal space region and is then nearest-neighbor interpolated to obtain isotropic sampling. The corresponding reconstructed diffraction pattern, obtained via Fourier transform of the electron density, is shown in 2(c). Comparison between 2(a) and 2(c) shows good agreement between the measured and reconstructed diffraction patterns, especially in brighter regions where the signal-to-noise ratio in the data is higher. In particular, the reconstructed pattern preserves the fringe spacing and overall contours of the data. It also features singularities (dark spots) in the already low-intensity regions of the diffraction pattern that are not clearly visible in the measured detector image. These characteristics persist across reconstructions with different random initial starting points. The apparent density of the reconstruction in 2(b) features a higher-density central region surrounded by a few-micron shell over which the density tapers off by more than an order of magnitude (Fig. 8). In an effort to understand whether the lower-magnitude outer regions seen in pale blue in 2(b) are physical or simply a reconstruction artifact, we perform a set of reconstructions with various support sizes in which the support is unaltered during the reconstruction algorithm. 2(d) compares a line cut of the data (black circles) to line cuts of the Fourier transform of the reconstructed density as a function of support diameter. The reconstructed reciprocal space intensities reproduce the features of the data well only for sufficiently large supports that extend well beyond the bounds of the higher-magnitude central region of the density. This analysis therefore suggests that there is indeed a physical basis for the lower-magnitude region featured in 2(b). Furthermore, we obtain similar results from data collected on a different SL in the same capillary (see Figs. 9, 10 and accompanying text). We defer proposing origins of this apparent lower-magnitude density region to the Discussion in order to first present the XPCS results for this same SL structure.
@@
To characterize aspects of SL annealing we performed additional analysis on the dataset shown in Fig. 3. To elucidate the scale-dependent dynamics within the SL, we calculate correlation functions as a function of distance from the Bragg peak, as shown in Fig. 4(a). We divide the detector ROI containing the peak into annular regions bounded by contours of constant , and we further subdivide each region along the longitudinal and transverse components of , marked in gold and blue, respectively. In this SAXS geometry, the longitudinal direction is parallel to , and the transverse direction is tangent to the (111) diffraction ring. For each annular region, we fit the corresponding autocorrelation function, , to a compressed/stretched exponential of the form , where is a baseline, is the speckle contrast, is the decorrelation rate, and is a stretching exponent. Fig. 4(b)-(c) shows example autocorrelations and fits for respective longitudinal and transverse directions at (see also Figs. 11, 12, 13 for data and fits at additional values). Such stretched/compressed exponentials often appear in systems with heterogeneous dynamics, where the resulting distribution of time scales leads to distortion of the simple exponential decay or Gaussian correlation forms [43]. Note that a speckle contrast higher than 1 arises when there are large variations in intensity (e.g., diffraction fringes) within the ROI. Because this analysis depends on the choice of center for the annular regions of interest and there is ambiguity in how this choice should be made, we repeat the analysis for many different choices of center, using pixels within a circular region (indicated in red in Fig. 4(a)) with a 5 pixel radius around the intensity center of mass, and averaging the resultant fit parameters together. Fig. 4(d) shows the decorrelation rate as a function of for each of the radial and transverse directions, respectively. The corresponding stretching exponents, , are shown in Fig. 4(e). The behavior of at low (region i, in gray) is complex and non-monotonic. The dispersion contrasts with that of, for instance, a simple, freely-diffusing colloidal system, where [26]. Unlike a diffusive system, which has no characteristic length scale, however, our system has a clear length scale of relevance, namely the size of the SL. For values around this value we would not anticipate a simple dispersion. But at sufficiently small length scales we should expect to recover monotonicity, which we indeed observe at high (region ii, white). Here the identification of the two regions is merely to highlight the point between these regimes at which such conventional, monotonic behavior manifests in the correlation functions. In the transverse direction, the decorrelation rate is on the order of Hz and increases with increasing , while in the longitudinal direction we do not resolve such dispersive behavior (and cannot reliably extract fits at the highest values). This trend in decorrelation rate in the transverse direction, and the lack thereof in the longitudinal direction, persist even if one fits to a fixed exponent of instead (Fig. 14). In order to investigate the source of this wavevector-dependent time scale, we calculate a two-time correlation function in Fig. 4(f) using the ring-shaped region of Fig. 4(a) around Two-time analysis enables us to observe from when in the data acquisition period the time scale arises and whether it is irreversible [44, 45, 46]. The raw data were first binned down to 0.1 s intervals to more readily visualize larger jumps in intensity. We observe distinct square-like patterns in the two-time correlation, corresponding to sudden shifts in the diffraction intensity at 17 s and 90 s.
@@ III Discussion
This work extends the reach of coherent diffractive imaging by probing a solid, immobilized structure in its native liquid environment [47]. To be able to extract meaningful conclusions about SL structure and evolution from our diffraction data via CDI and XPCS analysis techniques, the SLs must be sufficiently immobilized in the X-ray beam. The low- behavior of the SL Bragg peak is governed by long-range order in the SL; thus the oscillations that we observe in the autocorrelation functions in 3, which feature prominently at low , are likely produced by a bulk motion of the SL, such as small librations of the SL in solution. Perturbed by solvated particles or external vibrations, the SLs sway periodically on the quartz wool, producing intensity fluctuations at a rate much slower than the X-ray pulse repetition rate of 120 Hz. While the SLs are therefore not perfectly immobilized in the capillary, these perturbations are quite small (see Appendix F and Figs. 15, 16), and we are nevertheless able to resolve sharp reciprocal space features in our dataset, an encouraging sign that the sample is sufficiently immobilized for the purposes of our XPCS and CDI analyses.
@@
We now turn to a discussion of the long-time behavior of the SLs, as revealed by our XPCS results. We first observe that the dispersive increase (Fig. 4(d)) at high in the transverse fit parameters extracted from the XPCS data reflects conventional dynamical behavior in which scattering decorrelates more quickly at higher wavevector () [26]. Moreover, wavevectors in the transverse direction are sensitive to shear-like motions of the SL, analogous to the sensitivity to transverse acoustic phonons near Bragg peaks [48, 49]. Therefore, since we observe this dispersion along the transverse direction but not along the longitudinal direction, we interpret the time scale as that of shear-like fluctuations of the SL. We see, furthermore, that this one-time decorrelation is accompanied by square-like patterns in the two-time correlation (Fig. 4(f)), a phenomenon that we also observe in an additional experimental dataset (Fig. 17). In a third experimental dataset (Figs. 18, 19), we do not observe any decay in the autocorrelation functions at long delay, and the corresponding two-time correlation function (Fig. 20) is highly uniform and featureless, displaying no such rapid discontinuities as those seen in Fig. 4(f). Since the absence or presence of a one-time decorrelation rate is thus accompanied by the absence or presence of these discontinuities in the two-time correlation, the decorrelation is likely dominated by sudden changes in the diffraction pattern instead of small fluctuations therein. While we cannot rule out the possibility that these sudden changes in the diffraction pattern are merely due to bulk translations or rotations of the SL in solution, the increasing dispersion in decorrelation rate at high is indicative of internal SL dynamics. Furthermore, from prior ensemble SAXS measurements on NC SLs [13], we expect SLs to continue to evolve (via growth and defect annealing) even many hours after they are initially nucleated, provided their solution phase environment is maintained. Taken together, these features in the data suggest that the high- fluctuations in the SL Bragg peak are driven by sudden shear-like “avalanches” [50, 51] of NCs rather than a continuous annealing process. Such events are roughly analogous to the microstructural avalanches previously observed in martensitic metallic alloys [46, 45]. In our previous in situ SAXS experiments of SL formation, SLs were shown to evolve at different rates depending on the quench depth at which they were formed. At shallow quenches, highly ordered SLs formed but did not evolve much after their initial formation. In contrast, at slightly deeper quenches, more disordered SLs form initially but they continue to evolve over the course of the several hour observation period [13]. This observation is consistent with a picture in which annealing proceeds via such shear-like avalanches, in which these strain-relieving events are more likely to occur in regions of high disorder [52, 53, 54]. Moreover, previous ensemble measurements also find that the SL Bragg peak width evolves according to a power law over time, a relationship we would expect to arise from the occurrence of avalanches within many SLs within an ensemble throughout a bulk measurement window [55].
@@
Returning to the real space reconstruction obtained from CDI presented in Fig. 2, we consider interpretations of the decrease in electron density that arises moving outward from the center of the SL. One possibility is that the reconstruction algorithm has failed, assigning low non-zero values to regions of real space that extend beyond the boundary of the object. Another possibility is that the diffraction data is blurred due to the SL librations discussed above, thus yielding artifacts on large length scales in the reconstruction. A further alternative is that inhomogeneities in the beam path (due to the presence of solvent, quartz wool, and capillary glass) induce an effective decoherence in the beam and distort the diffraction signal. There are several reasons, however, to believe that the low-density regions are not an experimental artifact. First, the analysis of size-dependence presented in 2(d) suggests that there is diffracting material over a span of 2-3 m, extending beyond the high density region of the reconstruction. Additionally, we observe similar such low-density regions in the second CDI dataset that we analyzed (Fig. 10). For this latter dataset the reconstructed object, including the low-density region, is smaller, in accordance with a larger fringe spacing. If beam contamination were the driving force behind a reconstruction artifact, we would expect to see its impact appear on similar length scales across all datasets. Furthermore, we can estimate an upper bound for the magnitude of phase distortion in the scattered beam. For quartz wool fibers of 9-30 m, supposing that the part of the beam passes through at worst the edge of one fiber, we would expect a relative phase shift of around . This places an upper bound on the phase distortion due to the quartz wool, and is small enough to preserve a good CDI signal (see Appendix C for quantitative estimates). One physical interpretation of these apparent lower electron densities is that they are produced by regions of greater disorder in the SL. By analyzing and comparing hypothetical ordered and disordered SL structures, we attempt to assess the impact of at least one kind of defect on such a Bragg CDI reconstruction.
@@
The combination of CDI and XPCS analysis together has proven to be a valuable means to obtain a coherent description of a self-assembled SL structure and its evolution in solution that could not otherwise have been realized. We previously found via in situ SAXS that initial SL nucleation takes place on a rapid (sub-second) time scale, and then growth and annealing progress over the course of minutes and hours. Taking all of our findings together, one interpretation of the apparent density profile we observe from CDI is that the SL crystallinity decays away from its center, with regions of greater disorder towards the surface. While we cannot rule out this possibility based on the reconstructions alone, we would not expect such a disordered shell to be so large in proportion to the crystalline core. Although other phases such as gels and liquids have been implicated as products of NC self-assembly [1], we found no evidence in this work and in previous work with the same exact metal NC system [13] of these non-crystalline condensed phases. In addition, real space optical and electron microscopy images of this system following a quench show large, micron sized polycrystalline structures (Fig. 6). Therefore a more likely explanation is that the SL (111) peak analyzed in Fig. 2 is from a single crystal grain of a larger cluster of aggregated SLs, each of which nucleated separately; that is, the reconstruction “core” is simply one large, optimally oriented domain among many. Our analysis of hypothetical polycrystals in Fig. 5 demonstrates that in performing CDI on a polycrystalline structure whose other grains do not meet the Bragg condition, we would expect to see regions of apparent low density in the reconstructed object. This polycrystalline picture also accords well with the account of sub-lattice dynamics indicated by correlation analysis and with the structures observed in electron microscopy in Fig. 6(c)-(d). In particular, the presence of defects such as slip planes or grain boundaries between neighboring crystallites can lead to a build-up of strain in the SL. Once this strain becomes sufficiently large, annealing should occur via discrete, large-scale shear motions. Based on the data analyzed here and in previous work, we thus propose that Au NC SLs quickly nucleate and then coalesce into polycrystalline structures, with shear annealing occurring within and/or between crystallites via strain-relieving avalanches.